OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation. x2 + x - 12 = 0

OpenStudy (anonymous):

u got this one mon

OpenStudy (anonymous):

lol do i do it the same way as the others?

OpenStudy (anonymous):

you can just factor it normal way

OpenStudy (anonymous):

you dont need to use the quadratic formula

OpenStudy (anonymous):

then how do i factor it

OpenStudy (anonymous):

look for the x and the number ..

OpenStudy (anonymous):

you know you get the factors of 12 and they have to equal b when you add them

OpenStudy (jiteshmeghwal9):

x^2+4x-3x-12=0

OpenStudy (jiteshmeghwal9):

x(x+4)-3(x+4)=0 (x-3)(x+4)=0 hence,x=3 & -4